OpenStudy (anonymous):

EASY MEDAL!!! solve fully for a medal <333. The scale on a map is 1 inch = 40 miles. Round to the nearest mile. On the map it is 7.5 inches from Chicago to Cincinnati. Traveling at 65 mi/h, how long will it take to drive from Chicago to Cincinnati? Round to the nearest tenth of an hour. a. 4.6 hours c. 8.7 hours b. 5.2 hours d. 12.0 hours

OpenStudy (pradius):

4.6

OpenStudy (pradius):

O.o thanks

OpenStudy (anonymous):

7.5*40 then /by 65=4.61

OpenStudy (anonymous):

1. If the map says it is 18 1/8 inches from point A to point B, and 1 inch = 40 miles then you multiply 18 1/8 by 40. This is because each inch is the equivalent of 40 miles. For example, 1 inch is 40 miles, 2 inches is 80 miles (because 40 * 2 = 80), 3 inches would be 120 miles etc. 1/8 = 0.125 so 18 1/8 = 18.125 So just take a calculator (or just use google, that's always easy) and enter 18.125 * 40. Your answer should come out to exactly 725. 2. First you have to find how far 2.5 inches is. This is just like the last question. Just multiply 2.5 by 40, and you get 100. Therefore, it is 100 miles from Indianapolis to Terra Haute, IN. You know that it is 185 miles from Chicago to Indianapolis, so to find the total, just add them together. 185 + 100 = 285. 3. First, you got it, you have to find how far you are going. 7.5 * 40 = 300. So now you know you are traveling 300 miles. This part can get a little tricky at first, but once you get it, it's easy. It's the same concept as finding how many miles are in a given number of inches, except the units have different names. If you travel 65 miles, in 1 hour, then you travel 130 miles in 2 hours, and 195 miles in 3 hours etc. To find how many times 65 goes into 300, in other words, how many sets of 65 miles can you make if you are traveling a total of 300 miles, you have to divide 300 by 65. 300 / 65 = 4.61538462. The question asks you to round to the nearest tenth of an hour, so the third number (the number in the hundredths place) determines if you round up, or round down. I'm assuming that you know this already, but I'm going to go through it anyway just because I can. A 0-4 rounds down, and a 5-9 rounds up. The number that determines what you round to in this case is a 1, which means it rounds down to a 0. Therefore, your rounded answer is 4.6 hours.